OpenStudy (anonymous):

Currently, Sam's coach is twice as old as Sam. When Sam started soccer 10 years ago, his coach was four times his age. How old is Sam now?

OpenStudy (anonymous):

show me how to do it and also show me the expression for how to find coach's age 10 years ago yhank you

OpenStudy (anonymous):

c = 4s but not exactly

OpenStudy (anonymous):

i know sam age is 15

OpenStudy (anonymous):

vut i dont know coacch expression for 10 years ago

OpenStudy (anonymous):

*but

OpenStudy (anonymous):

well 10 years ago same was 5, so you can use that too

OpenStudy (anonymous):

which means 10 years ago sam's coach was 20

OpenStudy (anonymous):

but I need the expression for coach

OpenStudy (anonymous):

20 = 4(15) - c 20 = 60 - c -40 = -c => 40 = c so, 20 = 4s - 40

OpenStudy (anonymous):

20 = 4s - c

OpenStudy (anonymous):

let age of sam = x let age of his coach = y calculate the value of x by using following equation 2x = y 4(x-10) = y